Problem 6. Italia Espresso Machina Inc. produces a single product. Data concerning the companys operations last year appear below Units in beginning inventory Units produced Units sold 2,000 1,900 Selling price per unit $100 Variable costs per unit Direct materials Direct labor Variable manufacturing overhead Variable selling and administrative $30 $10 $5 $2 Fixed costs in total Fixed manufacturing overhead Fixed selling and administrative 40,000 $60,000 Required a. Compute the unit product cost under both absorption and variable costing b. Prepare an income statement for the year using absorption costing c. Prepare a contribution format income statement for the year using variable costing d. Prepare a report reconciling the difference in net operating income between absorption and variable costing for the year
Answer #1

Answers

  • All working forms part of the answer
  • Requirement ‘a’

Working

Absorption Costing

Variable Costing

Variable cost per unit:

Direct material

$                   30.00

$                 30.00

Direct Labor

$                   10.00

$                 10.00

Variable manufacturing overhead

$                     5.00

$                    5.00

A

Total Variable manufacturing cost per unit

$                   45.00

$                 45.00

B

Fixed manufacturing overhead

$          40,000.00

C

Units produced

                      2,000

D = B/C

Fixed manufacturing cost per unit

$                   20.00

E = A + D

Unit Product Cost

$                   65.00

$                 45.00

  • Requirement ‘b’

--Absorption costing income statement

A = 1900 units x $ 100

Sales Revenue

$             190,000

Cost of Goods Sold:

B

Beginning Inventory

$                          -  

C = 2000 units x $ 65

Cost of production

$              130,000

D = 100 units x $ 65

Ending Inventory

$                (6,500)

E = 1900 units x $ 65

Cost of Goods Sold

$             123,500

F = A - E

Gross Profit

$               66,500

Operating Expenses:

G = 1900 units x $ 2

Variable selling & administrative expenses

$                   3,800

H

Fixed Selling & administrative expense

$                60,000

I = G+H

Total Operating Expenses

$               63,800

J = F - I

Net Operating Income

$                 2,700

  • Requirement ‘c’

--variable Costing Income Statement

A = 1900 units x $ 100

Sales Revenue

$           190,000

Variable costs:

B = 1900 units x $30

Direct material

$         57,000

C = 1900 units x $10

Direct Labor

$         19,000

D = 1900 units x $5

variable manufacturing cost

$            9,500

E = 1900 units x $2

Variable selling & admin cost

$            3,800

F = B+C+D+E

Total Variable cost

$              89,300

G = A - F

Contribution margin

$           100,700

Fixed costs:

H

Fixed manufacturing cost

$         40,000

I

Fixed selling & administrative

$         60,000

J = H + I

Total Fixed cost

$           100,000

K = G - J

Net Operating Income

$                    700

  • Requirement ‘d’

--Under Absorption costing, Fixed manufacturing overhead of $ 20 per unit [see requirement 1] is deferred to next period in Ending Inventory.

Reconciliation:

A

Net Operating Income as per Absorption costing

$                   2,700

B = 100 units x $ 20

Less: Amount of Fixed manufacturing overhead deferred in Ending Inventory

$                   2,000

C = A - B

Net Operating Income as per variable Costing

$                      700